OpenStudy (anonymous):

(True or false.) Let V be an n-dimensional vector space, and let S be a k-dimensional subspace of V, where k < n. If {⃗v1,...,⃗vk} is a basis for S and {⃗vk+1,...,⃗vn} is a set of n − k linearly independent vectors not in S, then {⃗v1,...,⃗vk,⃗vk+1,...,⃗vn} is a basis for V.

OpenStudy (anonymous):

its false. but i don't understand why

OpenStudy (anonymous):

It is indeed false. Example: \(V=\mathbb{R}^3\) \(v_1=(1,0,0)\) \(S=\text{span}\ \{v_1\}\) \(v_2=(1,1,1)\) \(v_3=(0,1,1)\) Notice that \({v_2,v_3}\) is linearly independent and \(v_2,v_3 \not\in S\). Yet \({v_1,v_2,v_3}\) is not linearly independent since \(v_1 = v_2 − v_3\). Therefore \({v_1,v_2,v_3}\) is not a basis for \(\mathbb{R}^3\).

OpenStudy (anonymous):

Oh right! Thank you

OpenStudy (anonymous):

another question: Also, why doesn't the combined span of a subspace, say V in Rn and its orthogonal complement Vperp, do not span Rn. It says so in my lecture notes but doesn't prove why.

OpenStudy (anonymous):

I don't know what you mean by "combined span" but if \(V \subset \mathbb{R}^n\) is a vector subspace if \(\mathbb{R}^n\) then \(\displaystyle V \oplus V^\perp = \mathbb{R}^n\)

OpenStudy (anonymous):

from my lecture notes: "Suppose W a subspace of R . Show by an example that W ∪ W⊥ (The expression W ∪ W⊥ means taking W and W⊥ together) need not be the whole vector space R3 Solution: Simply consider the plane P containing the origin. We have seen that P⊥ is a line perpendicular to P containing 0. We see that P ∪ P is not all of R3.

OpenStudy (anonymous):

From my lecture notes: Suppose W a subspace of R3 Show that \[W \cup W ^{⊥}\] need not be the whole vector space R . Solution: Simply consider the plane P containing the origin. We have seen that P is a line perpendicular to P containing 0. We see that \[P \cup P ^{\perp}\] is not all of R3 .

OpenStudy (anonymous):

Yes, that makes sense. Do you understand why?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

P u P will be 3 LI vectors, right?

OpenStudy (anonymous):

and i thought 3 LI vectors can span R3

OpenStudy (anonymous):

You are not taking the span of the vectors in the two spaces. You are simply considering the union.

OpenStudy (anonymous):

Think of just gluing together the two spaces.

OpenStudy (anonymous):

Oh! ....

OpenStudy (anonymous):

Got it! thank you very much

OpenStudy (anonymous):

\(P \cup P^\perp\) is just the set of all vectors in \(P\) or \(P^\perp\), the union. I'm sure you know what that means.

OpenStudy (anonymous):

The union is not necessarily even a vector space or vector subspace.

OpenStudy (anonymous):

so, If there was a question, Let U={v1,v2} be a plane in R3 and U(perp)={v3} , the set, {v1,v2,v3} spans R3. This would be true, right?

OpenStudy (anonymous):

Yes if you consider the span then it is indeed \(\mathbb{R}^3\). However, let's look at a picture of the union.

OpenStudy (anonymous):